Circular No. 5919

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET KUSHIDA (1994a) Additional observations communicated by S. Nakano, Sumoto, Japan:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
Jan. 9.55382 9^{h}46^{m}20.^{s}05 +3^{o}03′08′′.3 12.5 Urata
9.70764 9 46 24.54 +3 01 34.1 12.5 Sugie
9.70903 9 46 24.41 +3 01 31.4 13  Koishikawa
9.72500 9 46 25.02 +3 01 23.6  Sugie
9.73090 9 46 25.07 +3 01 18.8  Koishikawa
10.80243 9 46 58.85 +2 50 39.1 12.0 Kushida
10.81632 9 46 59.18 +2 50 28.7  _{”}
11.65964 9 47 24.36 +2 42 30.2 13.2 Kawanishi
11.66211 9 47 24.42 +2 42 29.6  _{”}
11.70772 9 47 25.33 +2 42 03.9 12.5 Urata
11.71091 9 47 25.38 +2 42 02.0  _{”}
11.71253 9 47 25.42 +2 42 01.1  _{”}
11.71410 9 47 25.43 +2 42 00.2  _{”}

T. Urata (Nihondaira Obs.). Jan. 9.55382 observation is remeasurement of

one on IAUC 5918. A. Sugie (Dynic Astronomical Observatory). 0.25-m f/3.4 Schmidt. M. Koishikawa (Sendai Obs., Ayashi Station). 0.30-m f/3.8 reflector. Y. Kushida (Yatsugatake South Base Observatory). 0.25-m f/3.4 reflector. K. Kawanishi (Akou). 0.33-m f/3.3 reflector + CCD.

Nakano provides the following preliminary parabolic orbital elements from 25 observations, Jan. 9-11; this may be a short-period comet.

T = 1993 Dec. 5.327 TT  ω = 198.^{o}979
Ω = 254.335 }2000.0
q = 1.36087 AU  i = 4.968

SUPERNOVA 1993aj IN ANONYMOUS GALAXY M. Hamuy, Cerro Tololo Interamerican Observatory, reports the fol-

lowing magnitudes (± 0.10) of SN 1993aj, based on observations gathered by J. Storm: 1993 Dec. 29.333 UT, B = 18.58, V = 18.15; 1994 Jan. 2.320, 18.76, 18.22.

1994 January 11 Daniel W. E. Green